OpenStudy (anonymous):

If an = 3n + 8, generate the 16th term.

jimthompson5910 (jim_thompson5910):

Assuming that n starts at n=1, the 16th term is \[a_{16}=3(16)+8 =48+8 = 56\] So the 16th term is \[a_{16}=56\]